SERIES CIRCUIT:

In the circuit, the total voltage source is 30.0 V, R1 = 4.0 x 10^1 Ω and R2 = 15 Ω. After 10 seconds, how many coulombs of charge has passed through the circuit?
1
Expert's answer
2019-12-24T14:33:35-0500

The figure shows a series circuit with two resistances and a voltage source. Since the circuit is series through the resistance flows the same current "I". According to Ohm's law, the voltage drop across the resistances is "U_1=I\cdot R_1" ; "U_2=I\cdot R_2" According to Kirchhoff's voltage law, the sum of all voltages over any closed loop is . In our case, this means "U-U_1-U_2=0" . From the equation one can determine the magnitude of the current "U=I\cdot (R_1+R_2)";

"I=\frac{U}{R_1+R_2}=\frac{30.0V}{(40+15)\Omega}=0.55 A".

By definition, the current is equal to the charge transferred per unit time, i.e. "I=\frac{\Delta Q}{\Delta t}". Thus the charge equals "\Delta Q=I\cdot \Delta t=0.55 A\cdot 10s=5.5 C" .

Answer: After 10 seconds, "5.5 C" has passed through the circuit.